# Deflated Dynamics Value Iteration

### v1

@@ 8.  Proof of Theoretical Results
$=U\left((1-\alpha)I+\alpha\gamma R+e_{1:s}R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}e_{1:s}^{\top}\right)U^{\top}$
where $R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}$ is $s\times s$ upper triangular matrix with digonal entries $\frac{(\lambda_{1}\gamma-1)\gamma\alpha^{2}\lambda_{1}}{1-\gamma\alpha\lambda_{1}},\dots,\frac{(\lambda_{s}\gamma-1)\gamma\alpha^{2}\lambda_{s}}{1-\gamma\alpha\lambda_{s}}$, satisfying $R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}=D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}+\alpha\gamma(R_{s}-D_{s,\lambda_{i}})D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}$. By simple calculation, we know that $(1-\alpha)I+\alpha\gamma R+e_{1:s}R_{\frac{(\alpha\lambda_{i}\gamma-\alpha)\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}e_{1:s}^{\top}$ is $n\times n$ upper triangular matrix with digonal entries $\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{1}},\dots,\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{s}},(1-\alpha)+\alpha\gamma\lambda_{s+1},\dots,(1-\alpha)+\alpha\gamma\lambda_{n}$.
@@
8.3.  Proof of Theorem 4.1
By definition of DDVI for Control (3),
$V^{k}-V^{\star}=\left(I+\frac{\gamma}{1-\gamma}E_{1}\right)\left(W^{k}-\left(I-\gamma E_{1}\right)V^{\star}\right)$
@@
$W^{k}-W^{\star}=\gamma\left({\mathcal{P}^{\pi_{g}}-E_{1}}\right)W^{k-1}+r^{\pi_{g}}-\gamma\left({\mathcal{P}^{\pi^{\star}}-E_{1}}\right)W^{\star}-r^{\pi^{\star}}$
where $\pi_{g}=\operatorname*{argmax}_{\pi}\{r^{\pi}+\gamma({\mathcal{P}}^{\pi}-E_{1})W^{k-1}\}$ and $\pi^{\star}$ is optimal policy. This implies that
@@
$t_{i}\left({\gamma\mathcal{P}^{\pi^{\star}}(W^{k-1}-W^{\star})}\right)_{i}+(1-t_{i})\left({\gamma\mathcal{P}^{\pi_{g}}(W^{k-1}-W^{\star})}\right)_{i}=\left({W^{k}-W^{\star}+\gamma E_{1}(W^{k-1}-W^{\star})}\right)_{i}$
@@
and this implies that
@@
$\|V^{k}-V^{\star}\|\leq\frac{2}{1-\gamma}\gamma^{k}\|V^{0}-V^{\star}\|$
@@
Suppose there exist unique optimal policy $\pi^{\star}$. Then, if (non-deflated) VI generates $V^{k}$ for $k=0,1,\dots$, there exist $K$ such that $\operatorname*{argmax}_{\pi}T^{\pi}V^{k}=\pi^{\star}$ for $k>K$. Since DDVI for Control generates same policy with VI for Control, DDVI for Control also generates greedy policy $\pi^{\star}$ for $k>K$ iterations. Therefore, we get
@@
8.4.  Proof of Theorem 5.1
Consider following stochastic approximation algorithm
$Y^{k+1}=Y^{k}+\eta_{k}f(Y^{k},\zeta_{k})$ (5)
for $k=0,1,\dots$, where $Y^{k}\in^{n}$, $\eta_{k}\in^{+}$, $f(\cdot,z):^{n}\rightarrow^{n}$ is uniformly Lipschitz, and $\{\zeta_{k}\}_{k=0,1,\dots}$ are i.i.d random variables. Let $F(y)=\mathbf{E}[f(y,\zeta_{k})]$, $F_{\infty}(y)=\lim_{r\rightarrow\infty}F(ry)/r$, $M^{k+1}=f(Y^{k},\zeta_{k})-F(Y^{t})$, and $\mathcal{F}^{k}=\sigma(Y^{i},M^{i},1\leq i\leq t)$. Then following proposition holds.
Proposition 8.1.
[Borkar and Meyn, 2000, Theorem 2.2 and 2.5] If (i) $\eta_{k}=(k+1)^{-1}$, (ii) $\{M^{k},\mathcal{F}^{k}\}_{k=0,1,\dots,}$ are martingale difference sequence , (iii) $\mathbf{E}[M^{k+1}\,|\,\mathcal{F}^{k}]\leq K(1+\|Y^{k}\|^{2})$ for some constant $K$, (iv) $\dot{y}(t)=F_{\infty}(y(t))$ has asymptotically stable equilibrium origin, and (v) $\dot{y}(t)=F(y(t))$ has a unique globally asymtotically stable equilibrium $y^{\star}$, then $\{Y^{t}\}_{k=0,1,\dots}$ of (5) converges to $y^{\star}$ almost surely, and furthermore, $\{Y^{t}\}_{k=0,1,\dots}$ of
$Y^{k+1}(i_{k})=Y^{k}(i_{k})+\eta_{\nu_{i_{k},t}}f(Y^{k},\zeta_{k})(i_{k})$ (6)
for $k=0,1,\dots$ where $i_{k}$ is random variable taking a value in $\{1,2,\cdots,n\}$, $\nu_{i,t}=\sum^{k}_{k=0}\mathbf{1}_{\{i=i_{k}\}}$ satisfying lim $\liminf\nu_{i,t}/t>c$ for some constant $c>0$, also converge to $y^{\star}$ almost surely.
In (6), $Y^{k}(i_{k})$ and $f(Y^{k},\zeta_{k})(i_{k})$ are $i$-th coordinate of $Y^{k}$ and $f(Y^{k},\zeta_{k})$, respectively. (6) can be interpreted as asynchronous version of (5), and we note that Proposition 8.1 is a simplified version with stronger conditions of Theorem 2.2 and 2.5 of Borkar and Meyn [2000].
Recall that DDTD (4) with $\alpha=1$ is
$W^{k+1}(X_{k})$ $=W^{k}(X_{k})+\eta_{\nu_{i,t}}\big{[}r^{\pi}(X_{k})+\gamma V^{k}(X_{k}^{\prime})-\gamma(E_{s}V^{k})(X_{k})-W^{k}(X_{k})\big{]}$
@@
We will first show that $W^{k}\rightarrow(I-\gamma E_{s})V^{\pi}$ and this directly implies that $V^{k}\rightarrow V^{\pi}$. For applying Proposition 8.1, consider
$W^{k+1}=W^{k}+\eta_{k}f(W^{k},Z_{t})$
where $\eta_{k}=(k+1)^{-1}$, $Z_{t}\in^{n}$ such that $Z_{t}(x)\sim\mathcal{P}^{\pi}(\cdot\,|\,x)$ and $\{Z_{t}\}_{k=0,1,\dots}$ are i.i.d random variables, and $f(W^{k},Z_{t})(x)=r^{\pi}(x)+\gamma(\left(I-\gamma E_{s}\right)^{-1}W^{k})(Z_{t}(x))-\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}W^{k}(x)-W^{k}(x)$. Then, $F(W)=\mathbf{E}[f(W,Z_{t})]=[\gamma(\mathcal{P}^{\pi}-E_{s})(I-\gamma E_{s})^{-1}-I]W+r^{\pi}$, $M^{k+1}=f(W^{k},Z_{t})-F(W^{k})$, and $\mathcal{F}^{k}=\sigma(W^{i},M^{i},1\leq i\leq t)$.
We now check the conditions. First, since $f(W,z)(x)-f(W^{\prime},z)(x)=\left(I-\gamma E_{s}\right)^{-1}(W-W^{\prime})(z(x))-(I+\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1})(W-W^{\prime})(x)$ implies $\|f(W,z)-f(W^{\prime},z)\|\leq(\|\left(I-\gamma E_{s}\right)^{-1}\|+\|(I+\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1})\|)\|W-W^{\prime}\|$, $f(\cdot,z)$ is uniformly Lipschitz.
@@
for all $x\in\mathcal{X}$. This implies that $M^{k+1}$ is martingale difference sequence. Also, $\mathbf{E}[\|M^{k+1}\|^{2}\,|\,\mathcal{F}^{k}]\leq(\|\gamma\left(I-\gamma E_{s}\right)^{-1}\|+\|\gamma\mathcal{P}^{\pi}(I-\gamma E_{s})^{-1}\|)^{2}\|W^{k}\|^{2}\leq K(1+\|{W^{k}}\|^{2})$ for constant $K=(\|\gamma\left(I-\gamma E_{s}\right)^{-1}\|+\|\gamma\mathcal{P}^{\pi}(I-\gamma E_{s})^{-1}\|)^{2}$.
@@
By definition, $F_{\infty}(y)=\lim_{r\rightarrow\infty}F(ry)/r=\gamma(\mathcal{P}^{\pi}-E_{s}))(I-\gamma E_{s})^{-1}-I]y$. In Section 8.1, we showed that $\text{spec}(F_{\infty}+I)=\{\gamma\lambda_{s+1},\dots,\gamma\lambda_{n},0\}$. Hence, $\text{spec}(F_{\infty})=\{\gamma\lambda_{s+1}-1,\dots,\gamma\lambda_{1}-1,-1\}$. Since $|\lambda_{i}|\leq 1$ for all $s+1\leq i\leq n$, real parts of all eigenvalues $F_{\infty}$ are negative. This implies that $\dot{y}(t)=F_{\infty}(y(t))$, has an asymtotically stable equilibrium origin and $\dot{y}(t)=F(y(t))$ has a unique globally asymtotically stable equilibrium $W^{\pi}=[I-(\gamma(\mathcal{P}^{\pi}-E_{s}))(I-\gamma E_{s})^{-1}]^{-1}r^{\pi}$.
@@
Lastly, since $X_{k}\sim\text{Unif}(\mathcal{X})$ and $\{X_{k}\}_{k=0,1,\dots}$ are i.i.d random variables, $\nu_{i,t}/t$ converges to $\frac{1}{n}$ almost suerly by law of large numbers. Since $\eta_{\nu_{X_{k},t}}=(\sum^{k}_{i=0}\mathbf{1}_{X_{k}=X_{i}})^{-1}$, Therefore, by Proposition 8.1, $\{W^{k}\}_{k=0,1,\dots}$ of (7) converge to $W^{\pi}$ almost surely, and this implies that $V^{k}\rightarrow(I-\gamma E_{s})^{-1}W^{\pi}=(I-\gamma\mathcal{P}^{\pi})^{-1}r^{\pi}=V^{\pi}$ almost surely.
8.5.  Proof of Theorem 5.2
Consider following linear stochastic approximation algorithm
$Y^{k+1}=Y^{k}+\eta_{k}(A(\zeta_{k})Y^{t}+b(\zeta_{k}))$ (8)
for $k=0,1,\dots$, where $Y^{k}\in^{n}$, $\{\zeta_{k}\}_{k=0,1,\dots,}$ are random variables, $b(\zeta_{k})\in^{n}$, $A(\zeta_{k})\in^{n}\times^{n}$, and $\|A_{i,j}\|,\|b_{k}\|<\infty$, for $1\leq i,j\leq n$ and $1\leq k\leq n$. Then, following Proposition holds.
Proposition 8.2.
@@
[Chen et al., 2020, Theorem 2.5 and 2.7] If (i) $\eta_{k}=g(k+1)^{-1}$ for $g>0$, (ii) $\{\zeta_{k}\}_{k=0,1,\dots}$ is ergodic (aperiodic and irreducible) Markov process with unique invariant measure $\mu$, (iii) $\mathbf{E}_{\mu}[A^{t}]=A$ and $\mathbf{E}_{\mu}[B^{t}]=b$, (iv) $A$ is Hurwitz, (v) $Re(\lambda)<-1/2$ for all $\lambda\in$ spec($gA$), then $\mathbf{E}[\|Y^{t}-Y^{\star}\|^{2}]=O(k^{-1})$ where $Y^{\star}=-A^{-1}b$.
@@
We note that Proposition 8.1 is also a simplified version with stronger conditions of Theorem 2.5 and 2.7 of Chen et al. [2020].
@@
Define $\psi(X_{k})\in^{n}$ as $\psi(X_{k})(x)=\mathbf{1}_{\{X_{k}=x\}}$. Then, DDTD (4) with $\alpha=1$ is equivalent to
$W^{k+1}=W^{k}+\eta_{k}(A(X_{k},X^{\prime}_{k})W^{k}+b(X_{k},X^{\prime}_{k})),$
where $A(X,X^{\prime})=\psi(X)(\psi(X^{\prime})^{\top}\gamma\left(I-\gamma E_{s}\right)^{-1}-\psi(X)^{\top}\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}-\psi(X)^{\top})$ and $b(X,X^{\prime})=\psi(X)r^{\pi}(X)$. Hence $\|A_{i,j}\|,\|b_{k}\|<\infty$, for $1\leq i,j\leq n$ and $1\leq k\leq n$.
@@
Since $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ are i.i.d random variables, $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ are Markov process. Let $\mu\sim\{X_{k},X^{\prime}_{k}\}$ and $\mathcal{P}(x_{1},x^{\prime}_{1}\,|\,x_{2},x^{\prime}_{2})$ be transition matrix of $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$. Then $\mathcal{P}(x_{1},x^{\prime}_{1}\,|\,x_{2},x^{\prime}_{2})=\mu(x_{1},x_{1}^{\prime})=\frac{1}{n}\mathcal{P}^{\pi}(x^{\prime}_{1}\,|\,x_{1})$. Since $(\mu^{\prime})^{\top}\mathcal{P}(\cdot,\cdot\,|\,\cdot,\cdot)=\mu$ for any distribution $\mu^{\prime}$, $\mu$ is unique invarinat measure. Hence, $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ is irreducible and aperiodic.
@@
$=n^{-1}((\mathcal{P}^{\pi}-\gamma E_{s})\left(I-\gamma E_{s}\right)^{-1}-I)$
@@
As we discussed in Section 8.4, $\text{spec}(A)=\{\gamma\lambda_{s+1}-1,\dots,\gamma\lambda_{n}-1,-1\}$. Let $\lambda_{\text{DDTD}}=\min_{\lambda\in\{\lambda_{s+1},\dots,\lambda_{n}\}}\text{Re}(1-\gamma\lambda)$. Then, if $g>\frac{n}{2\lambda_{\text{DDTD}}}$, $Re(\lambda)<-1/2$ for all $\lambda\in$ spec($gA$). Therefore, by Proposition 8.2, $\mathbf{E}[\|W^{k}-W^{\pi}\|^{2}]=O(k^{-1})$ and this implies that $\mathbf{E}[\|V^{k}-V^{\pi}\|^{2}]\leq\mathbf{E}[\|(I-\gamma E_{s})^{-1}\|^{2}\|W^{k}-W^{\star}\|^{2}]=O(k^{-1})$.
8.6.  Non-asymptotic convergence analysis of DDTD

### v2

@@ Appendix B Proof of Theoretical Results
$=U\left((1-\alpha)I+\alpha\gamma R+e_{1:s}R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}e_{1:s}^{\top}\right)U^{\top},$
where $R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}$ is $s\times s$ upper triangular matrix with diagonal entries $\frac{(\lambda_{1}\gamma-1)\gamma\alpha^{2}\lambda_{1}}{1-\gamma\alpha\lambda_{1}},\dots,\frac{(\lambda_{s}\gamma-1)\gamma\alpha^{2}\lambda_{s}}{1-\gamma\alpha\lambda_{s}}$, satisfying $R_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}=D_{s,\frac{(\lambda_{i}\gamma-1)\gamma\alpha^{2}\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}+\alpha\gamma(R_{s}-D_{s,\lambda_{i}})D_{s,\frac{\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}$. By simple calculation, we know that $(1-\alpha)I+\alpha\gamma R+e_{1:s}R_{\frac{(\alpha\lambda_{i}\gamma-\alpha)\gamma\alpha\lambda_{i}}{1-\gamma\alpha\lambda_{i}}}e_{1:s}^{\top}$ is $n\times n$ upper triangular matrix with diagonal entries $\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{1}},\dots,\frac{(1-\alpha)}{1-\alpha\gamma\lambda_{s}},(1-\alpha)+\alpha\gamma\lambda_{s+1},\dots,(1-\alpha)+\alpha\gamma\lambda_{n}$.
@@
B.3 Proof of Theorem 4.1
By the definition of DDVI for Control (equation 6 and equation 7),
$V^{k}-V^{\star}=\left(I+\frac{\gamma}{1-\gamma}E_{1}\right)\left(W^{k}-\left(I-\gamma E_{1}\right)V^{\star}\right),$
@@
$W^{k}-W^{\star}=\gamma\left({\mathcal{P}^{\pi_{g}}-E_{1}}\right)W^{k-1}+r^{\pi_{g}}-\gamma\left({\mathcal{P}^{\pi^{\star}}-E_{1}}\right)W^{\star}-r^{\pi^{\star}},$
where $\pi_{g}=\operatorname*{arg\,max}_{\pi}\{r^{\pi}+\gamma({\mathcal{P}}^{\pi}-E_{1})W^{k-1}\}$ and $\pi^{\star}$ is an optimal policy. Then, by the definition of greedy policy, we have
$\gamma\left({\mathcal{P}^{\pi_{g}}-E_{1}}\right)W^{k-1}+r^{\pi_{g}}-\gamma\left({\mathcal{P}^{\pi^{\star}}-E_{1}}\right)W^{\star}-r^{\pi^{\star}}$ $\leq\gamma\left({\mathcal{P}^{\pi_{g}}-E_{1}}\right)(W^{k-1}-W^{\star}),$
$\gamma\left({\mathcal{P}^{\pi_{g}}-E_{1}}\right)W^{k-1}+r^{\pi_{g}}-\gamma\left({\mathcal{P}^{\pi^{\star}}-E_{1}}\right)W^{\star}-r^{\pi^{\star}}$ $\geq\gamma\left({\mathcal{P}^{\pi^{\star}}-E_{1}}\right)(W^{k-1}-W^{\star}).$
This implies that
@@
$t_{i}\left({\gamma\mathcal{P}^{\pi^{\star}}(W^{k-1}-W^{\star})}\right)_{i}+(1-t_{i})\left({\gamma\mathcal{P}^{\pi_{g}}(W^{k-1}-W^{\star})}\right)_{i}=\left({W^{k}-W^{\star}+\gamma E_{1}(W^{k-1}-W^{\star})}\right)_{i},$
@@
$=\gamma^{k}\prod_{i=1}^{k}\left({\mathcal{P}^{\pi_{i}}-E_{1}}\right)(V^{0}-V^{\star})$
@@
where the third and last equality follows from $(\mathcal{P}^{\pi_{i}}-E_{1})E_{1}=0$. This implies that
@@
$\|V^{k}-V^{\star}\|\leq\frac{2}{1-\gamma}\gamma^{k}\|V^{0}-V^{\star}\|,$
@@
Suppose there exists a unique optimal policy $\pi^{\star}$. Then, if (non-deflated) VI generates $V^{k}$ for $k=0,1,\dots$, there exist $K$ such that $\operatorname*{arg\,max}_{\pi}T^{\pi}V^{k}=\pi^{\star}$ for $k>K$. Since DDVI for Control generates the same policy as VI for Control, DDVI for Control also generates greedy policy $\pi^{\star}$ for $k>K$ iterations. Therefore, we get
@@
B.4 Proof of Theorem 5.1
Consider the following stochastic approximation algorithm
$Y^{k+1}=Y^{k}+\eta_{k}(X_{k})(X_{k})f(Y^{k},\zeta_{k})$ (11)
for $k=0,1,\dots$, where $Y^{k}\in^{n}$, $\eta_{k}(X_{k})\in^{+}$, $f(\cdot,z):^{n}\rightarrow^{n}$ is uniformly Lipschitz, and $\{\zeta_{k}\}_{k=0,1,\dots}$ are i.i.d. random variables. Let $F(y)=\mathbf{E}[f(y,\zeta_{k})]$, $F_{\infty}(y)=\lim_{r\rightarrow\infty}F(ry)/r$, $M^{k+1}=f(Y^{k},\zeta_{k})-F(Y^{t})$, and $\mathcal{F}^{k}=\sigma(Y^{i},M^{i},1\leq i\leq t)$. The following result from Borkar & Meyn (2000) shows the convergence of the $Y^{k}$ sequence.
Proposition B.1.
(Borkar & Meyn, 2000, Theorem 2.2 and 2.5) If (i) $\eta_{k}=(k+1)^{-1}$, (ii) $\{M^{k},\mathcal{F}^{k}\}_{k=0,1,\dots,}$ are martingale difference sequence , (iii) $\mathbf{E}[M^{k+1}\,|\,\mathcal{F}^{k}]\leq K(1+\|Y^{k}\|^{2})$ for some constant $K$, (iv) $\dot{y}(t)=F_{\infty}(y(t))$ has asymptotically stable equilibrium origin, and (v) $\dot{y}(t)=F(y(t))$ has a unique globally asymptotically stable equilibrium $y^{\star}$, then $\{Y^{t}\}_{k=0,1,\dots}$ of equation 11 converges to $y^{\star}$ almost surely, and furthermore, $\{Y^{t}\}_{k=0,1,\dots}$ of
$Y^{k+1}(i_{k})=Y^{k}(i_{k})+\eta_{\nu_{i_{k},k}}f(Y^{k},\zeta_{k})(i_{k})$ (12)
for $k=0,1,\dots$, where $i_{k}$ are random variables taking values in $\{1,2,\cdots,n\}$ and $\nu_{i_{k},k}=\sum^{k}_{t=0}\mathbf{1}_{\{i=i_{t}\}}$ when $i_{k}=i$ satisfying lim $\liminf\nu_{i_{k},k}/k>c$ for some constant $c>0$, also converge to $y^{\star}$ almost surely.
In equation 12, $Y^{k}(i_{k})$ and $f(Y^{k},\zeta_{k})(i_{k})$ are the $i$-th coordinate of $Y^{k}$ and $f(Y^{k},\zeta_{k})$, respectively. The equation 12 can be interpreted as asynchronous version of equation 11. We note that Proposition B.1 is a simplified version with stronger conditions of Theorem 2.2 and 2.5 of Borkar & Meyn (2000).
Recall that DDTD equation 8 with $\alpha=1$ is
$W^{k+1}(X_{k})$ $=W^{k}(X_{k})+\eta_{k}(X_{k})\big{[}r^{\pi}(X_{k})+\gamma V^{k}(X_{k}^{\prime})-\gamma(E_{s}V^{k})(X_{k})-W^{k}(X_{k})\big{]}$
@@
We will first show that $W^{k}\rightarrow(I-\gamma E_{s})V^{\pi}$ and this directly implies that $V^{k}\rightarrow V^{\pi}$. For applying Proposition B.1, consider
$W^{k+1}=W^{k}+\eta_{k}f(W^{k},Z_{t}),$
where $\eta_{k}=(k+1)^{-1}$, $Z_{t}\in^{n}$ such that $Z_{t}(x)\sim\mathcal{P}^{\pi}(\cdot\,|\,x)$ and $\{Z_{t}\}_{k=0,1,\dots}$ are i.i.d. random variables, and $f(W^{k},Z_{t})(x)=r^{\pi}(x)+\gamma(\left(I-\gamma E_{s}\right)^{-1}W^{k})(Z_{t}(x))-\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}W^{k}(x)-W^{k}(x)$. Then,
$F(W)=\mathbf{E}[f(W,Z_{t})]=[\gamma(\mathcal{P}^{\pi}-E_{s})(I-\gamma E_{s})^{-1}-I]W+r^{\pi},$
$M^{k+1}=f(W^{k},Z_{t})-F(W^{k})$, and $\mathcal{F}^{k}=\sigma(W^{i},M^{i},1\leq i\leq t)$.
We now check the conditions. First, since $f(W,z)(x)-f(W^{\prime},z)(x)=\left(I-\gamma E_{s}\right)^{-1}(W-W^{\prime})(z(x))-(I+\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1})(W-W^{\prime})(x)$ implies $\|f(W,z)-f(W^{\prime},z)\|\leq(\|\left(I-\gamma E_{s}\right)^{-1}\|+\|(I+\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1})\|)\|W-W^{\prime}\|$, $f(\cdot,z)$ is uniformly Lipschitz. We have
@@
for all $x\in\mathcal{X}$. This implies that $M^{k+1}$ is a martingale difference sequence. Also,
$\mathbf{E}[\|M^{k+1}\|^{2}\,|\,\mathcal{F}^{k}]\leq(\|\gamma\left(I-\gamma E_{s}\right)^{-1}\|+\|\gamma\mathcal{P}^{\pi}(I-\gamma E_{s})^{-1}\|)^{2}\|W^{k}\|^{2}\leq K(1+\|{W^{k}}\|^{2})$
for constant $K=(\|\gamma\left(I-\gamma E_{s}\right)^{-1}\|+\|\gamma\mathcal{P}^{\pi}(I-\gamma E_{s})^{-1}\|)^{2}$.
@@
By definition, $F_{\infty}(y)=\lim_{r\rightarrow\infty}F(ry)/r=\gamma(\mathcal{P}^{\pi}-E_{s}))(I-\gamma E_{s})^{-1}-I]y$. In Section B.1, we showed that $\text{spec}(F_{\infty}+I)=\{\gamma\lambda_{s+1},\dots,\gamma\lambda_{n},0\}$. Hence, $\text{spec}(F_{\infty})=\{\gamma\lambda_{s+1}-1,\dots,\gamma\lambda_{1}-1,-1\}$. Since $|\lambda_{i}|\leq 1$ for all $s+1\leq i\leq n$, real parts of all eigenvalues $F_{\infty}$ are negative. This implies that $\dot{y}(t)=F_{\infty}(y(t))$ has an asymptotically stable equilibrium origin and $\dot{y}(t)=F(y(t))$ has a unique globally asymptotically stable equilibrium $W^{\pi}=[I-(\gamma(\mathcal{P}^{\pi}-E_{s}))(I-\gamma E_{s})^{-1}]^{-1}r^{\pi}$.
@@
Lastly, since $X_{k}\sim\text{Unif}(\mathcal{X})$ and $\{X_{k}\}_{k=0,1,\dots}$ are i.i.d. random variables, $\nu_{i,t}/t$ converges to $\frac{1}{n}$ almost surely by the law of large numbers. Since $\eta_{\nu_{X_{k},k}}=(\sum^{k}_{i=0}\mathbf{1}_{X_{k}=X_{i}})^{-1}$, Therefore, by Proposition B.1, $\{W^{k}\}_{k=0,1,\dots}$ of equation 13 converge to $W^{\pi}$ almost surely, and this implies that $V^{k}\rightarrow(I-\gamma E_{s})^{-1}W^{\pi}=(I-\gamma\mathcal{P}^{\pi})^{-1}r^{\pi}=V^{\pi}$ almost surely.
B.5 Proof of Theorem 5.2
Consider the following linear stochastic approximation algorithm
$Y^{k+1}=Y^{k}+\eta_{k}(X_{k})(A(\zeta_{k})Y^{t}+b(\zeta_{k}))$ (14)
for $k=0,1,\dots$, where $Y^{k}\in^{n}$, $\{\zeta_{k}\}_{k=0,1,\dots,}$ are random variables, $b(\zeta_{k})\in^{n}$, $A(\zeta_{k})\in^{n}\times^{n}$, and $\|A_{i,j}\|,\|b_{k}\|<\infty$, for $1\leq i,j\leq n$ and $1\leq k\leq n$. The following result from Chen et al. (2020) provides a rate of convergence of $Y^{k}$ to the solution $Y^{\star}=-A^{-1}b$, where $A$ and $b$ are the expectation of matrices $A(\zeta_{k})$ and vectors $b(\zeta_{k})$, respectively.
Proposition B.2.
@@
(Chen et al., 2020, Theorem 2.5 and Corollary 2.7) If (i) $\eta_{k}=g(k+1)^{-1}$ for $g>0$, (ii) $\{\zeta_{k}\}_{k=0,1,\dots}$ is ergodic (aperiodic and irreducible) Markov process with a unique invariant measure $\mu$, (iii) $\mathbf{E}_{\mu}[A^{t}]=A$ and $\mathbf{E}_{\mu}[B^{t}]=b$, (iv) $A$ is Hurwitz, (v) $Re(\lambda)<-1/2$ for all $\lambda\in$ spec($gA$), then $\mathbf{E}[\|Y^{t}-Y^{\star}\|^{2}]=O(k^{-1})$ where $Y^{\star}=-A^{-1}b$.
@@
We note that Proposition B.2 is also a simplified version with stronger conditions of Theorem 2.5 and Corollary 2.7 of Chen et al. (2020).
@@
Define $\psi(X_{k})\in^{n}$ as $\psi(X_{k})(x)=\mathbf{1}_{\{X_{k}=x\}}$. Then, DDTD equation 8 with $\alpha=1$ is equivalent to
$W^{k+1}=W^{k}+\eta_{k}(X_{k})(A(X_{k},X^{\prime}_{k})W^{k}+b(X_{k},X^{\prime}_{k})),$
where
$A(X,X^{\prime})=\psi(X)(\psi(X^{\prime})^{\top}\gamma\left(I-\gamma E_{s}\right)^{-1}-\psi(X)^{\top}\gamma E_{s}\left(I-\gamma E_{s}\right)^{-1}-\psi(X)^{\top})$
and $b(X,X^{\prime})=\psi(X)r^{\pi}(X)$. Hence $\|A_{i,j}\|,\|b_{k}\|<\infty$, for $1\leq i,j\leq n$ and $1\leq k\leq n$.
@@
Since $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ are i.i.d. random variables, $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ are Markov process. Let $\mu\sim\{X_{k},X^{\prime}_{k}\}$ and $\mathcal{P}(x_{1},x^{\prime}_{1}\,|\,x_{2},x^{\prime}_{2})$ be transition matrix of $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$. Then $\mathcal{P}(x_{1},x^{\prime}_{1}\,|\,x_{2},x^{\prime}_{2})=\mu(x_{1},x_{1}^{\prime})=\frac{1}{n}\mathcal{P}^{\pi}(x^{\prime}_{1}\,|\,x_{1})$. Since $(\mu^{\prime})^{\top}\mathcal{P}(\cdot,\cdot\,|\,\cdot,\cdot)=\mu$ for any distribution $\mu^{\prime}$, $\mu$ is a unique invariant measure. Hence, $\{X_{k},X^{\prime}_{k}\}_{k=0,1,\dots}$ is irreducible and aperiodic. We have,
@@
$=n^{-1}((\mathcal{P}^{\pi}-\gamma E_{s})\left(I-\gamma E_{s}\right)^{-1}-I),$
@@
As we discussed in Section B.4, $\text{spec}(A)=\{\gamma\lambda_{s+1}-1,\dots,\gamma\lambda_{n}-1,-1\}$. Let $\lambda_{\text{DDTD}}=\min_{\lambda\in\{\lambda_{s+1},\dots,\lambda_{n}\}}\text{Re}(1-\gamma\lambda)$. Then, if $g>\frac{n}{2\lambda_{\text{DDTD}}}$, the real component of the eigenvalues $Re(\lambda)<-1/2$ for all $\lambda\in$ spec($gA$). Therefore, by Proposition B.2, $\mathbf{E}[\|W^{k}-W^{\pi}\|^{2}]=O(k^{-1})$ and this implies that $\mathbf{E}[\|V^{k}-V^{\pi}\|^{2}]\leq\mathbf{E}[\|(I-\gamma E_{s})^{-1}\|^{2}\|W^{k}-W^{\star}\|^{2}]=O(k^{-1})$.
B.6 Non-asymptotic convergence analysis of DDTD